# GreenScan: Towards large-scale monitoring the health of urban trees using mobile sensing

### 22 Dec 2023 v1

@@ III-A2 Software Modules
2. Image Registration: Image registration involves matching or aligning images taken by two different sensors into a single coordinate system for further analysis [41]. It includes detecting key points from one image and mapping them to another image. Since both the multispectral and thermal imaging sensors have different FOVs (field of view) and are un-aligned, this modules aligns the multispectral images to the thermal images through linear translation in both horizontal and vertical directions. Also, to compensate for wider FOV of the multispectral sensor, this module also handles zooming in on the multispectral images.
@@
3. Image Segmentation: The is the most computationally intensive software module of the system. Recall that the aim of our system is to calculate the NDVI and CTD values for each tree in the images. However, these values should be calculated only for the leaves in tree canopy excluding the wooden parts, such as the trunk and branches. This is solved using a fusion of custom developed Mask R-CNN and pixel-wise NDVI analysis. Given a multispectral RGN image captured using the multispectral imaging sensor, this task can be broken into two sub-problems as follows:
  • Detect the canopy part of the trees even in cases where the image contains multiple trees: This is solved using a custom-developed Mask R-CNN model. The Mask R-CNN model is trained using transfer learning and discussed in more detail in Section III-B. It segments the instances of the tree canopies in the RGN image by generating a mask (segmentation) over them as shown in Figure 6.
@@
With the MAPIR Survey 3W employed in the prototype, a value of 0.02 was used as the cutoff to eliminate non-vegetation elements in the image. This value was derived using the analysis of the images captured during data collection experiments. An instance of inputs and outputs utilizing this module is also shown in Figure 4.
@@
The CTD value is computed by calculating the raw temperature value for each pixel by converting its color intensity value in the grayscale thermal image as per (2), computing the mean temperature over all pixels in the canopy and subtracting the ambient air temperature from the mean canopy temperature as per (3). CTD is calculated as:
@@
where $NIR$ and $Red$ are values of near-infrared channel and visible red channel for each pixel respectively. Our corrected NDVI is calculated as:
@@
III-B Development of Custom Mask R-CNN
@@
For the system to operate autonomously, the images will be captured in an unsupervised fashion. Thus, in addition to multiple trees in a single image, they may contain other objects such as cars, buildings, grass, and snow. Hence, it is imperative to individually identify all the tree canopies in an image and feed them to the calculation and analysis module. The custom mask R-CNN part of the image segmentaion module solves this by providing instance segmentation of the tree canopies in the image. To our knowledge, there is no pre-existing model available for instance segmentation of tree canopies or even trees in standard RGB images. The problem is further complicated as our input is RGN (Red, Green, Near-infrared) images from the multispectral imaging sensor instead of standard RGB images. For instance, we found that pre-trained models like Deeplabv3 [44], which can perform semantic segmentation of trees and vegetation on standard RGB images, perform poorly on RGN images.
III-B1 Training Data
@@
Any deep learning model requires training data in order to optimise the weights and activations of the layers. However, there does not exist a dataset with labels for instances of trees or tree canopies for RGN Images. Hence, we manually created the dataset using the RGN images collected during the data collection experiments (See Section IV-B). Each tree canopy in the image was manually annotated using a popular image annotation tool called LabelMe [45]. During annotation, only tree canopies that were completely present in the image were labelled. After this process, our dataset consisted of 51 annotated RGN images with two classes namely tree canopies and background.
III-B2 Training Process and Training Curve
@@
Our dataset consists of a relatively small number of images to train a deep learning model like Mask R-CNN from scratch. Transfer learning combined with data augmentation was employed in order to develop a custom model by using an existing model pre-trained on a different dataset. For this, we used a Mask R-CNN pre-trained [46] on COCO[47] (a dataset with 330K images) with ResNet101 as the backbone. We re-trained only the head layers (the top layers without the backbone) on our dataset. The batch size was configured as 4 and the number of epochs was 10. The training was performed on the Google Cloud Platform with a N1 instance with 13GB memory and 2vCPUs. We also generated synthetic data by augmenting the original dataset with flips in the horizontal and vertical directions and applying Gaussian blur. This increased the training dataset size by 50% and acted as a regularizer. The manually annotated dataset (refer Section III-B1) consisting of 51 images was split in the ratio of 70: 30 for training: testing. During re-training, each epoch took approximately 3 hours on the N1 instance. The training curve of the model is shown in Figure 5. It is seen from the training curve that only a small number of epochs are sufficient to reach the optimal validation loss on the test set owing to the retraining of only the head layers. The visual output results from our model are shown in Figure 6.
@@
Figure 5: The training curve of Mask R-CNN with epochs=10 and batch size=4, the red point indicates point of minimum loss, training losses as defined in [48]
@@
(b) Segmentation output from our Custom Mask R-CNN (instance segmentation model) trained using transfer learning
@@ III-B3 Model Quantization
Mask R-CNN is a relatively heavy model both from training and inference points of view. Hence, the developed Mask R-CNN was optimized to run on the edge at the cost of possible minute performance reduction. For this, the model built on TensorFlow was converted to TensorFlow-lite with dynamic range quantization [49]. Dynamic range quantization means that only the weights of the layers in 32-bit Floats in the full model are stored as 8-bit INTs while the activations of the layers are quantized during runtime. Our custom-made Mask R-CNN built over TensorFlow took around 15 seconds per inference of an image on a Raspberry Pi 4 while the TensorFlow-lite model reduced the inference time to 7 seconds with one-fourth of the CPU usage as the original TensorFlow model.
IV Evaluation
@@
We evaluated our system using a ground truth dataset obtained from the municipality of Cambridge, USA. We also conducted three data collection experiments to measure attributes of urban trees using GreenScan. In this section, we elaborate on these datasets followed by the obtained results.
IV-A (Ground truth) Tree Health Dataset
@@
Municipalities in cities obtain ground truth tree health data through city-wide surveys over years. For instance, in the city of Cambridge, USA, a survey is performed every 5 years whereas, for the city of Delft, The Netherlands, a survey is performed every 2 years. For the evaluation of our work, we obtained a ground truth tree health dataset for the city of Cambridge, USA through the Cambridge Urban Forest Master Plan [14]. A 2018 dataset was obtained through a combination of manual in-person arborist visits, satellite-based remote sensing and aerial LiDAR [14]. The dataset classifies the health conditions of trees into three categories, namely good, poor, and fair. The dataset contains information about 47,063 trees out of which 35,821 are in good health, 5176 are in fair health and 6066 are in poor health. Hence, most of the trees ($>75\%$) are rated as having a good health condition. In addition, the dataset contains information about the tree species, common name, the satellite-based NDVI, the latitude and the longitude, location, the shape length and shape area of the canopy, flood tolerance, and drought tolerance. This dataset was provided as Shapefiles (.shp, a data format used by Geographical Information Systems (GIS)) and was loaded to the online platform CARTO [50] (a GIS and spatial analysis tool). The staleness of data in terms of time further necessitates the advancement in this field of tree health monitoring.
IV-B Data Collection Experiments
@@
We collected multispectral (RGN) and thermal images through the developed system on three separate days in Cambridge, USA during the month of February 2022. A push button was used as the event trigger for the system. Hence, we employed the developed GreenScan system as a citizen science project with the 3D printed casing (pedestrian moving at walking speed in a straight line at a distance of 8m - 20m from the tree). In total, we collected data for 49 trees spread over two species namely Red Pine and Eastern White Pine trees. The multispectral imaging sensor was configured with a shutter speed of 1/60s and ISO of 50. The thermal imaging sensor was configured to measure temperature in range of (-10, 40) ${}^{\circ}$. The sites of data collection experiments are shown in Figure 7.
@@
Data Cleaning: During the first day of the data collection experiments, the Raspberry Pi hung up due to unknown reasons leading to a forced restart. On the third day of the experiments, owing to cold temperatures, the power supply had to be changed during data collection. These interruptions and restarts resulted in unstable values for a sequence of readings related to the canopy temperature by the thermal imaging sensor. As a result, these 11 data points were removed from our dataset generated using data collection experiments. In the end, our dataset was reduced to contain 40 trees. Distribution of the data collected from each of the tree species after data cleaning is shown in Table III.
@@
TABLE III: Distribution of trees after data cleaning
IV-C Performance of custom Mask R-CNN
To measure the performance of our custom Mask R-CNN model, we calculated the standard evaluation metrics [51] as used by COCO. Specifically, we measured mean Average Precision (mAP) / Average Precision (as per [51]) at different IoU (Intersection over Union) thresholds. The performance of our custom Mask R-CNN with and without quantization is shown in Table IV. A comparison of inference time and model size comparing both the full model and the quantized model is also shown in Table IV. In order to measure the stability of our results, k-Fold cross-validation was also performed with k=3, to evaluate the performance of the model on different training and test splits as shown in Table V. These results for different train-test splits as shown in Table V showcase the reliability of our model. From Table IV, it is seen that there is no significant reduction in performance using quantization. The inference time of the quantized model is half compared to the non-quantized model along with the reduction of the model size.
An example of segmentation outputs generated by the full model and quantized model on the same image is also shown in Figure 8. This also showcases the similar performance for both the full and the optimised model in a visual form.
From Table , it may appear that the $AP^{(IoU=0.5:0.95:0.05)}$ for the quantized model is increased slightly compared to the full model. On further exploring this anomaly, it was found that this behaviour is exhibited due to our annotated dataset where most images contain only one full tree canopy as ground truth. Thus, a model (non-quantized model) generalising better to find partially visible tree canopies, in addition to the full tree canopy is penalised in terms of Precision (False Positive). Further, it is seen from Figure 9 that the performance of the quantized model decreases more than the full model at higher IoUs (IoU= 0.85 for quantized model compared to 0.90 for full model) signifying that it is slightly poorer at object localisation compared to the full model.
@@ IV-C Performance of custom Mask R-CNN
TABLE IV: Performance of custom R-CNN model (Full and Quantized model)
@@
TABLE V: Results of 3-Fold Cross Validation of custom Mask R-CNN model
@@
Figure 8: Outputs from custom Mask R-CNNs in Tensorflow and Tensorflow-lite
Figure 9: The AP scores with increasing IoU thresholds as per COCO metrics [47] for the full and quantized tf-lite model model
IV-D Results for the health of trees
@@
Figure 10: Variation of measured NDVI vs Remote NDVI for trees observed during data collection experiments
Pearson’s correlation coefficient (r) was measured to calculate the strength of the linear relationship between our measured values and ground truth data. The correlation matrix comprising all of our measured values with the three ground truth parameters namely Ground Truth Condition, Remote NDVI, and Area is shown in Figure 11. Further, the correlation results between the measured NDVI and CTD with ground truth parameters is shown in Table VI. From Table VI, it is seen that there is a moderately strong correlation (r=0.54 with $p<0.05$) between our measured NDVI and remote NDVI. For context, in recent works [52], the correlation between NDVI measured using two different satellites was found to be 0.74. Two datasets can be highly correlated but strongly disagree. Hence, a Bland-Altman plot [53] widely used to showcase the agreement between two measurement methods (instead of the strength of relationship as indicated by pearson’s coefficient) is shown in Figure 12. From the Bland-Altman plot, it is seen that there is strong agreement between the two methods (Remote NDVI and Terrestrial NDVI) with 98% (more than 95%) points lying between average difference ± 1.96 standard deviation of the difference limits.
@@
Figure 12: The Bland-Altman plot showcasing the agreement between Corrected NDVI and Remote NDVI. The dashed-middle line shows the mean difference. The top most and bottom most lines indicate the upper and lower lines of agreement respectively.
@@ 0.28 Yes
TABLE VI: The correlation between our measured values and ground truth parameters
The distribution of CTD and NDVI with respect to ground truth health conditions is shown in Figure 13. From the NDVI distribution in Figure 12(a), it is seen that the extent of agreement of NDVI with respect to the ground truth health conditions varies with respect to the species. For red-pine trees, the trees in good health condiiton have higher measured NDVI values than trees in poor and fair condition. A similar conclusion is drawn from the CTD distribution in Figure 12(b). The mean NDVI and CTD for each species is also shown in Table VII.
@@
TABLE VII: The mean of measured NDVI and CTD across species and health
IV-D1 High-level tree health analysis
From Figure 11, it is clear that there is almost no correlation between NDVI and CTD. Thus, they are independently measuring two different attributes related to tree health and useful to incorporate in the system. Further, in recent works such as [52], the correlation between remote NDVI measured using two different satellites was found to be 0.74 (moderately strong). From Table VI, the moderately strong correlation (r=0.54 with $p<0.05$) between our measured NDVI (ground based) and remote NDVI can be mainly due to the time drift (four years apart) between the measurements. In any case, this moderately strong correlation serves to showcase the validity of our approach. The weak-moderate correlation (r=0.28 with $p<0.05$) between CTD and ground truth tree health condition can be attributed to the skewed distribution of the dataset where more trees are rated as having good conditions compared to poor and fair conditions. Further analysis of CTD distribution for all trees in Figure 12(b) shows the high variability of CTD for trees in poor condition leading to this overall weak-moderate correlation. Interestingly, due to unknown reasons, NDVI is found to have a significant and weak moderate correlation (r=0.28 with $p<0.05$) with the area of tree canopy obtained from the ground truth dataset in Table VI.
IV-D2 Species-wise tree health analysis
It is inferred from Table VII and Figure 13, that the distribution of NDVI values are in agreement with the health condition of trees for both red pine and eastern white pine species. In fact, mean NDVI values for Red pine as shown in Table VII are significant ($p<0.05$) for good and poor condition trees. The mean NDVI values for Eastern White Pine trees are insignificant ($p>0.05$ considering the three different tree health conditions). However, from the NDVI distributions for Eastern White pine in Figure 12(a), it is seen that the good condition trees are generally distributed to have higher NDVI values than poor and fair condition trees.

### 06 Apr 2024 v2

@@ 3.1.2 Software Modules
2. Image Registration: Image registration involves matching or aligning images taken by two different sensors into a single coordinate system for further analysis [49]. It includes detecting key points from one image and mapping them to another image. Since both the multispectral and thermal imaging sensors have different FoVs (field of views) and are un-aligned, this modules aligns the multispectral images to the thermal images through linear translation in both horizontal and vertical directions. Also, to compensate for wider FOV of the multispectral sensor, this module also handles zooming in on the multispectral images.
@@
3. Image Segmentation: The is the most computationally intensive software module of the system. Recall that the aim of our system is to calculate the NDVI and CTD values for each tree in the images. However, these values should be calculated only for the leaves in tree canopy excluding the wooden parts, such as the trunk and branches. This is solved using a fusion of custom developed Mask R-CNN (Mask Regional Convolutional Neural Network) and pixel-wise NDVI analysis. Mask R-CNN [51] is an object detection and instance segmentation model that identifies and then draws a precise mask around the detected object. Given a multispectral RGN image captured using the multispectral imaging sensor, this task can be broken into two sub-problems as follows:
  • Detect the canopy part of the trees even in cases where the image contains multiple trees: This is solved using a custom-developed Mask R-CNN model. The Mask R-CNN model is trained using transfer learning and discussed in more detail in Section 3.2. It segments the instances of the tree canopies in the RGN image by generating a mask (segmentation) over them as shown in Figure 5.
@@
With the MAPIR Survey 3W employed in the GreenScan system, a value of 0.02 was used as the cutoff to eliminate non-vegetation elements in the image. This value was derived using the analysis of the images captured during data collection experiments. An instance of inputs and outputs utilizing this module is also shown in Figure 3.
@@
The CTD value is computed by calculating the raw temperature value for each pixel in the grayscale thermal image as per (2), computing the mean temperature over all pixels in the canopy and subtracting the ambient air temperature from the mean canopy temperature as per (3). CTD is calculated as:
@@
where $NIR$ and $Red$ are values of near-infrared channel and visible red channel for each pixel respectively. The corrected NDVI is calculated as:
@@
3.2 Development of Custom Mask R-CNN
@@
For the system to operate autonomously, the images will be captured in an unsupervised fashion. Thus, in addition to multiple trees in a single image, they may contain other objects such as cars, buildings, grass, and snow. Hence, it is imperative to individually identify all the tree canopies in an image and feed them to the Analysis and Calculation Module. The custom mask R-CNN part of the image segmentation module solves this by providing instance segmentation of the tree canopies in the image. To our knowledge, there is no pre-existing model available for instance segmentation of tree canopies or even trees in standard RGB images. The problem is further complicated as our input is RGN (Red, Green, Near-infrared) images from the multispectral imaging sensor instead of standard RGB images. For instance, we found that pre-trained models like Deeplabv3 [53], which can perform semantic segmentation of trees and vegetation on standard RGB images, perform poorly on RGN images.
3.2.1 Training Data
@@
Any deep learning model requires training data in order to optimise the weights and activations of the layers. However, there does not exist a dataset with labels for instances of trees or tree canopies for RGN Images. Hence, we manually created the dataset using the RGN images collected during the data collection experiments (See Section 4.2). Each tree canopy in the image was manually annotated using a popular image annotation tool called LabelMe [54]. During annotation, only tree canopies that were completely present in the image were labelled. After this process, our dataset consisted of 51 annotated RGN images with two classes namely tree canopies and background.
3.2.2 Training Process and Training Curve
@@
Our dataset consists of a relatively small number of images to train a deep learning model like Mask R-CNN from scratch. Transfer learning combined with data augmentation was employed in order to develop a custom model by using an existing model pre-trained on a different dataset. For this, we used a Mask R-CNN pre-trained [55] on COCO [56] (a dataset with 330K images) with ResNet101 as the backbone. We re-trained only the head layers (the top layers without the backbone) on our dataset. The batch size was configured as 4 and the number of epochs was 10. The training was performed on the Google Cloud Platform with a N1 instance with 13GB memory and 2vCPUs. We also generated synthetic data by augmenting the original dataset with flips in the horizontal and vertical directions and applying Gaussian blur. This increased the training dataset size by 50% and acted as a regularizer. The manually annotated dataset (refer Section 3.2.1) consisting of 51 images was split in the ratio of 70: 30 for training: testing. During re-training, each epoch took approximately 3 hours on the N1 instance. The training curve of the model is shown in Figure 4. It is seen from the training curve that only a small number of epochs are sufficient to reach the optimal validation loss on the test set owing to the retraining of only the head layers. The visual output results from our model are shown in Figure 5.
@@
Figure 4: The training curve of Mask R-CNN with epochs=10 and batch size=4. The red point indicates point of minimum loss and training losses are as defined in [51]
@@
(b) Segmentation output from our Custom Mask R-CNN instance segmentation model trained using transfer learning
@@ 3.2.3 Model Quantization
Mask R-CNN is a relatively heavy model both from training and inference points of view. Hence, the developed Mask R-CNN was optimized to run on the edge at the cost of possible minute performance reduction. For this, the model built on TensorFlow was converted to TensorFlow-lite with dynamic range quantization [57]. Dynamic range quantization means that only the weights of the layers in 32-bit FLOAT of the full model are stored as 8-bit INTs while the activations of the layers are quantized during runtime. Our custom-made Mask R-CNN built over TensorFlow took around 15 seconds per inference of an image on a Raspberry Pi 3 while the TensorFlow-lite model reduced the inference time to 7 seconds with one-fourth of the CPU usage as the original TensorFlow model.
4 Evaluation
@@
The system was evaluated using a dataset obtained from the municipality of Cambridge, USA as ground truth reference. We also conducted three data collection experiments to collect data of urban trees using the GreenScan system. In this section, we elaborate on this dataset and the data collection experiments followed by the obtained results.
4.1 (Ground truth) Tree Health Dataset
@@
Municipalities in cities obtain tree health data through city-wide surveys over years. For instance, in the city of Cambridge, USA, a survey is performed every 5 years whereas, for the city of Delft, The Netherlands, a survey is performed every 1 year to 2 years depending on the previously rated condition of the tree. For the evaluation of our work, we obtained the tree health dataset for the city of Cambridge, USA through the Cambridge Urban Forest Master Plan [16] to be used as ground truth reference. A 2018 dataset was obtained from the municipality. This dataset was created through a combination of manual in-person arborist visits, satellite-based remote sensing and aerial LiDAR [16]. The dataset classifies the health conditions of trees into three categories, namely good, poor, and fair. The dataset contains information about 47,063 trees out of which 35,821 are in good health, 5176 are in fair health and 6066 are in poor health. Hence, most of the trees ($>75\%$) are rated as having a good health condition. In addition, the dataset contains information about the tree species, common name, the satellite-based NDVI, the latitude and the longitude, location, the shape length and shape area of the canopy and other parameters. This dataset was provided as Shapefiles (.shp, a data format used by GIS (Geographical Information Systems)) and was loaded to the online platform CARTO [58] (a GIS and spatial analysis tool). On a side note, the staleness of data in terms of time also necessitates the advancement in this field of tree health monitoring.
4.2 Data Collection Experiments
@@
We collected multispectral (RGN) and thermal images through the developed system on three separate days in Cambridge, USA during the month of February 2022. A push button was used as the event trigger for the system. Hence, we employed the developed GreenScan system as a citizen science project with the 3D printed casing (pedestrian moving at walking speed in a straight line at a distance of 8m - 20m from the tree). In total, we collected data for 49 trees spread over two species namely Red Pine and Eastern White Pine trees. The multispectral imaging sensor was configured with a shutter speed of 1/60s and ISO of 50. The thermal imaging sensor was configured to measure temperature in range of (-10, 40) ${}^{\circ}$. The sites of data collection experiments are shown in Figure 6, chosen based on the species and accessibility.
@@
Data Cleaning: During the first day of the data collection experiments, the Raspberry Pi hung up due to unknown reasons leading to a forced restart. On the third day of the experiments, owing to cold temperatures, the power supply had to be changed during data collection. These interruptions and restarts resulted in unstable values for a sequence of readings related to the canopy temperature by the thermal imaging sensor. As a result, these 11 data points were removed from our dataset generated using data collection experiments. In the end, our dataset was reduced to contain 40 trees. Distribution of the data collected from each of the tree species after data cleaning is shown in Table 3.
Table 3: Distribution of trees after data cleaning
@@
4.3 Performance of custom Mask R-CNN
To measure the performance of our custom Mask R-CNN model, we calculated the standard evaluation metrics as used by COCO [59]. Specifically, we measured $mAP$ (mean Average Precision) / $AP$ (Average Precision) at different IoU (Intersection over Union) thresholds (as per [59]). The performance of our custom Mask R-CNN with and without quantization is shown in Table 4. A comparison of inference time and model size comparing both the full model and the quantized model is also shown in Table 4. In order to measure the stability of our results, k-Fold cross-validation [60] was also performed with k=3, to evaluate the performance of the model on different training and test splits as shown in Table 5. These results for different train-test splits as shown in Table 5 showcase the reliability of our model.
From Table 4, it is seen that there is no significant reduction in performance using quantization. The inference time of the quantized model is half compared to the non-quantized model along with the reduction of the model size. An example of segmentation outputs generated by the full model and quantized model on the same image is also shown in Figure 7. This also showcases the similar performance for both the full and the optimised model in a visual form.
From Table 4, it may appear that the $AP^{(IoU=0.5:0.95:0.05)}$ [59] for the quantized model is increased slightly compared to the full model. On further exploring this anomaly, it was found that this behaviour is exhibited due to our annotated dataset where most images contain only one full tree canopy as ground truth. Thus, a model (non-quantized model) generalising better to find partially visible tree canopies, in addition to the full tree canopy is penalised in terms of Precision (False Positive). Further, it is seen from Figure 8 that the performance of the quantized model decreases more than the full model at higher IoUs (IoU= 0.85 for quantized model compared to 0.90 for full model) signifying that it is slightly poorer at object localisation compared to the full model.
Table 4: Performance of custom R-CNN model (Full and Quantized model)
@@ 4.3 Performance of custom Mask R-CNN
Table 5: Results of 3-Fold Cross Validation of custom Mask R-CNN model
@@
Figure 7: Outputs from the full and quantized custom Mask R-CNNs
Figure 8: The AP scores with increasing IoU thresholds as per COCO metrics [56] for the full and quantized models
4.4 Results for the health of trees
@@
Figure 9: Variation of measured NDVI vs Remote NDVI for trees observed during data collection experiments. The tree index refers to an individual tree ID.
Two datasets can be highly correlated but strongly disagree. Hence, a Bland-Altman plot [61] widely used to showcase the agreement between two monitoring methods measuring the same attribute was utilised. It plots the difference between the corresponding measurement values against the average of those values. From the Bland-Altman plot shown in Figure 10, it is seen that there is strong agreement between the two methods (Remote NDVI and our measured NDVI) with all points (representing data for each tree) except one lying within the 95% limits of agreement (average difference within ± 1.96 standard deviation of the difference, as the difference between values follows a normal distribution).
Figure 10: The Bland-Altman plot showcasing the agreement between the our measured NDVI and Remote NDVI. The dashed-middle line shows the mean difference. The top most and bottom most lines refer to the 95% limits of agreement respectively.
Pearson’s correlation coefficient (r) was also measured to calculate the strength of the linear relationship between our measured values and ground truth data. The correlation matrix comprising all of our measured values with the three ground truth parameters namely Ground Truth Condition, Remote NDVI, and Area is shown in Figure 11. Further, the correlation results between the measured NDVI and CTD with ground truth parameters is shown in Table 6.
@@
Table 6: The correlation between our measured values and ground truth parameters
@@ 0.28 Yes
The distribution of CTD and NDVI with respect to ground truth health conditions is shown in Figure 12. From the NDVI distribution in Figure 11(a), it is seen that the extent of agreement of NDVI with respect to the ground truth health conditions varies with respect to the species. For red-pine trees, the trees in good health condition have higher measured NDVI values than trees in poor and fair condition. A similar conclusion is drawn from the CTD distribution in Figure 11(b). The mean NDVI and CTD for each species is also shown in Table 7.
@@
Table 7: The mean of measured NDVI and CTD across species and health
@@
4.4.1 High-level tree health analysis
From Figure 11, it is clear that there is almost no correlation between NDVI and CTD. Thus, they are independently measuring two different attributes related to tree health and useful to incorporate in the system. In recent works such as [62], the correlation between remote NDVI measured using two different satellites was found to be 0.74 (moderately strong). In this work, the Table 6 shows moderately strong correlation (r=0.54 with $p<0.05$) between our measured NDVI (ground based) and remote NDVI. This moderately strong correlation serves to showcase the validity of our approach for ground-based NDVI measurement using multispectral imaging sensors.
Since there is no ground truth reference attribute for CTD which indicates water stress of trees, we checked the correlation of CTD with ground truth health condition as shown in Table 6. The weak-moderate correlation (r=0.28 with $p<0.05$) between CTD and ground truth tree health condition can be attributed to the skewed distribution of the dataset where more trees are rated as having good conditions compared to poor and fair conditions. Further analysis of CTD distribution for all trees in Figure 11(b) shows the high variability of CTD for trees in poor condition leading to this overall weak-moderate correlation.
4.4.2 Species-wise tree health analysis
From the NDVI distributions for Eastern White pine in Figure 11(a), it is seen that the good condition trees are generally distributed to have higher NDVI values than poor and fair condition trees. Thus, a simple threshold based classification algorithm can easily flag trees which might not specifically be in good health conditions. At a scale of tens of thousands of trees in a city, this can lead to a significant amount of cost savings.